### 03 Dec 2023 v1

# A Cup Product Obstruction to Frobenius Stability

Abstract.
A countable discrete group is said to be Frobenius stable if a function from the group that is “almost multiplicative” in the point Frobenius norm topology is “close” to a genuine unitary representation in the same topology. The purpose of this note is to show that if $\Gamma$ is a finitely generated group and a non-torsion element of $H^{2}(\Gamma;\mathbb{Z})$ can be written as a cup product of two elements in $H^{1}(\Gamma;\mathbb{Z})$ then $\Gamma$ is not Frobenius stable. A nontrivial corollary is that Thompson’s Group $F$ is not Frobenius stable. Because the argument is the same we also show this result for stability in unnormalized Schatten $p$-norms for $1<p\leq\infty$.
1. Introduction
Let $\Gamma$ be a countable discrete group and let $1\leq p\leq\infty$. Let $||\cdot||_{p}$ denote the unnormalized Schatten $p$-norm, $||M||_{p}=(\operatorname{Tr}((M^{*}M)^{p/2}))^{1/p}$ for $p<\infty$ and operator norm for $p=\infty$. Then a discrete group $\Gamma$ is stable in the unnormalized Schatten $p$-norm if for all sequences of functions $\varphi_{n}:\Gamma\rightarrow U_{n}$ the condition
(1) $||\rho_{n}(gh)-\rho_{n}(g)\rho_{n}(h)||_{p}\rightarrow 0,\,\,\forall g,h\in\Gamma$
implies there exists a sequence of group homomorphisms $\varphi_{n}:\Gamma\rightarrow U_{n}$ so that
(2) $||\rho_{n}(g)-\varphi_{n}(g)||_{p}\rightarrow 0,\,\,\forall g\in\Gamma.$
Of particular interest is the $p=2$ case, called Frobenius stability, and the $p=\infty$ case called matricial stability. We will call a sequence of functions $(\varphi_{n})$ that satisfies condition (1) an asymptotic homomorphism. If there exist homomorphisms $(\psi_{n})$ satisfying condition (2) we say that $(\varphi_{n})$ is perturbable to homomorphisms. Frobenius stability was introduced by de Chiffre, Glebsky, Lubotzky, and Thom in [4].
The purpose of this note is to show the following observation.
Theorem 1.1.
Let $\Gamma$ be a finitely generated discrete group, and let $1<p\leq\infty$. If $\alpha,\beta\in H^{1}(\Gamma;\mathbb{Z})$ so that $\alpha\smile\beta\in H^{2}(\Gamma;\mathbb{Z})$ is non-torsion then $\Gamma$ is not stable in the unnormalized Schatten $p$-norm.
In the $p=\infty$ case our result follows from the methods developed by Dadarlat in [3], though it does not follow from the main result as stated there. The techniques we use here are more similar to those developed by the author in [7] and [8]. In [8] the notion of a skinny cohomology class is used as an obstruction to Frobenius stability of nilpotent groups. Here a cohomology class $[\sigma]$ is skinny with respect to a homomorphism $\alpha:\Gamma\rightarrow\mathbb{Z}$ if the restriction of $[\sigma]$ to the kernel of $\alpha$ is a coboundary. Then $\alpha\smile\beta$ is skinny with respect to both $\alpha$ and $\beta$, motivating our main result.
Corollary 1.2.
For $1<p\leq\infty$ Thompson’s group $F$ is not stable in the unnormalized Schatten $p$-norm.
Proof.
The cohomology ring of $F$ can be computed with methods in [11]; see [2] for an explicit computation in the case of $F$. From this computation, one can see that $H^{1}(F;\mathbb{Z})\cong\mathbb{Z}^{2}$ and the cup product of the two generators is non-torsion. ∎
This is an interesting contrast to the result of Fournier-Facio and Rangarajan in [6] that $F$ is uniformly stable with respect to all submultiplicative norms. Here by uniformly stable, we mean that the convergence in conditions (1) and (2) is uniform rather than pointwise.
The basic idea of the proof is that since $\alpha,\beta\in H^{1}(\Gamma;\mathbb{Z})\cong\operatorname{Hom}(\Gamma,\mathbb{Z})$ we can view the pair $(\alpha,\beta)$ as a homomorphism from $\Gamma$ to $\mathbb{Z}^{2}$. We use a classic example of an asymptotic homomorphism of $\mathbb{Z}^{2}$ due to Voiculescu in [12] and pull it back by $(\alpha,\beta)$ to get an asymptotic homomorphism of $\Gamma$. We show that this is a projective representation [^1] Meaning a map from $\Gamma$ to unitaries whose failure to be multiplicative, $\rho(gh)\rho(h)^{-1}\rho(g)^{-1}$, is scalar-valued. of $\Gamma$ (Lemma 3.1). Then we use a winding number type argument based on the nontriviality of $\alpha\smile\beta$ to show that this approximate representation is not perturbable (Lemma 3.2).
The asymptotic representation that is not perturbable can be described as follows.
Definition 1.3.
Since $H^{1}(\Gamma;\mathbb{Z})\cong\operatorname{Hom}(\Gamma,\mathbb{Z})$ we may view $\alpha$ and $\beta$ as group homomorphisms from $\Gamma$ to $\mathbb{Z}$. Then define
$\rho_{n}(g)=u_{n}^{\alpha(g)}v_{n}^{\beta(g)}$
where $u_{n}$ and $v_{n}$ are the $n\times n$ Voiculescu unitaries
$u_{n}=\begin{bmatrix}0&0&\cdots&0&0&1\\ 1&0&\cdots&0&0&0\\ 0&1&\cdots&0&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots&\vdots\\ 0&0&\cdots&1&0&0\\ 0&0&\cdots&0&1&0\end{bmatrix}\mbox{ and }v_{n}=\begin{bmatrix}\exp\left(\frac{2\pi i}{n}\right)&0&0&\cdots&0\\ 0&\exp\left(\frac{4\pi i}{n}\right)&0&\cdots&0\\ 0&0&\exp\left(\frac{6\pi i}{n}\right)&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\cdots&1\end{bmatrix}.$
2. Notation
There are many ways to characterize group homology and cohomology; we will give a concrete description of 1-cohomology, 2-cohomology, and 2-homology here. We will only use homology and cohomology with coefficients in $\mathbb{Z}$ and the trivial action in this paper. For more about this construction see [1, Chapter II.3].
As stated in the introduction, $H^{1}(\Gamma;\mathbb{Z})\cong\operatorname{Hom}(\Gamma,\mathbb{Z})$ and we can take this to be the definition.
Definition 2.1.
We define a 2-cocycle to be a function $\sigma$ from $\Gamma^{2}$ to $\mathbb{Z}$ satisfying the following equation
$\sigma(g,h)-\sigma(g,hk)+\sigma(gh,k)-\sigma(h,k)=0.$
A 2-coboundary is a function that can be written in the form
$\sigma(g,h)=\gamma(g)-\gamma(gh)+\gamma(h)$
for some function $\gamma:\Gamma\rightarrow\mathbb{Z}$. Every 2-coboundary is a 2-cocycle and $H^{2}(\Gamma;\mathbb{Z})$ is defined to be the group of 2-cocycles, mod the subgroup of 2-coboundaries. The group operation is pointwise addition.
Definition 2.2.
Call $C_{k}(\Gamma)$ to be formal linear combinations of elements of $\Gamma^{k}$. We write a typical element of $C_{2}(\Gamma)$ as
$\sum_{j=1}^{N}x_{j}[a_{j}|b_{j}]$
with $a_{j},b_{j}\in\Gamma$ and $x_{j}\in\mathbb{Z}$. Define $\partial_{2}:C_{2}(\Gamma)\rightarrow C_{1}(\Gamma)$ to by the equation
$\partial_{2}[a|b]=[a]-[ab]+[b]$
and $\partial_{3}:C_{3}(\Gamma)\rightarrow C_{2}(\Gamma)$ by
$\partial_{3}[a|b|c]=[a|b]-[a|bc]+[ab|c]-[b|c].$
Then $H_{2}(\Gamma;\mathbb{Z}):=\ker(\partial_{2})/\operatorname{im}(\partial_{3})$. An element of $\ker(\partial_{2})$ is referred to as a 2-cycle and an element in $\operatorname{im}(\partial_{3})$ is referred to as a 2-boundary.
Definition 2.3.
The Kronecker pairing between 2-homology and 2-cohomology is a bilinear map from $H^{2}(\Gamma;\mathbb{Z})\times H_{2}(\Gamma;\mathbb{Z})$ to $\mathbb{Z}$ defined by the formula
$\left\langle\sigma,\sum_{j=1}^{N}x_{n}[a_{j}|b_{j}]\right\rangle:=\sum_{j=1}^{N}x_{j}\sigma(a_{j},b_{j})$
where $\sigma$ is a cocycle, and $\sum_{j=1}^{N}x_{n}[a_{j}|b_{j}]$ is a cycle. The value does not depend on either choice of representative.
Definition 2.4.
[…]
$\sigma(g,h)=\alpha(g)\beta(h).$
For a more general definition, and more information see [1, Chapter V.3].
Definition 2.5 ([7] Definition 3.3).
If
$c=\sum_{j=1}^{N}x_{j}[a_{j}|b_{j}]\in C_{2}(\Gamma)$
and $\rho:\Gamma\rightarrow\operatorname{GL}_{n}(\mathbb{C})$ so that for all $j\in\{1,\ldots,N\}$
$||\rho(a_{j}b_{j})\rho(a_{j})^{-1}\rho(b_{j})^{-1}-\operatorname{id}_{\mathbb{C}^{n}}||<1$
we define
$\langle\rho,c\rangle=\frac{1}{2\pi i}\sum_{j=1}^{N}x_{j}\operatorname{Tr}(\log(\rho(a_{j}b_{j})\rho(b_{j})^{-1}\rho(a_{j})^{-1}))$
where $\log$ is defined as a power series centered at 1.
If $\partial c=0$ we have that $\langle\rho,c\rangle\in\mathbb{Z}$ [7, Proposition 3.4]. The version of the “winding number argument” we are using is as follows.
Theorem 2.6.
[7, Theorem 3.7] If $\rho_{0}$ is a (not necessarily unitary) representation of $\Gamma$, $\rho_{1}$ is a function from $\Gamma$ to $U(n)$,
$c=\sum_{j=1}^{N}x_{j}[a_{j}|b_{j}]$
is a 2-cycle on $\Gamma$, and
$||\rho_{1}(g)-\rho_{0}(g)||<\frac{1}{24}$
for all $g\in\{a_{j},b_{j},a_{j}b_{j}\}_{j=1}^{N}$ then $\rho_{1}$ is multiplicative enough for $\langle\rho_{1},c\rangle$ to be defined and $\langle\rho_{1},c\rangle=0$.
This argument has its roots in the “winding number argument” discovered by Kazhdan [9], and later independently by Exel and Loring [5].
3. Proofs
Lemma 3.1.
Define $\rho_{n}$ as in Definition 1.3. It obeys the following identity
$\rho_{n}(gh)\rho_{n}(h)^{-1}\rho_{n}(g)^{-1}=\exp\left(-\frac{2\pi i}{n}\beta(g)\alpha(h)\right)\operatorname{id}_{\mathbb{C}^{n}}.$
Proof.
Let $\omega=\exp\left(\frac{2\pi i}{n}\right)$ and note the four identities $v_{n}u_{n}=\omega u_{n}v_{n}$, $v_{n}^{-1}u_{n}=\omega^{-1}u_{n}v_{n}^{-1}$, $v_{n}u_{n}^{-1}=\omega^{-1}u_{n}^{-1}v_{n}$, and $v_{n}^{-1}u_{n}^{-1}=\omega u_{n}^{-1}v_{n}^{-1}$. From these four identities, it follows that $v_{n}^{x}u_{n}^{y}=\omega^{xy}u_{n}^{x}v_{n}^{y}$, for all $x,y\in\mathbb{Z}$. We compute
$\rho_{n}(gh)\rho_{n}(h)^{-1}\rho_{n}(g)^{-1}$ $=u_{n}^{\alpha(g)+\alpha(h)}v_{n}^{\beta(g)+\beta(h)}v_{n}^{-\beta(h)}u_{n}^{-\alpha(h)}v_{n}^{-\beta(g)}u_{n}^{-\alpha(g)}$
$=u_{n}^{\alpha(g)+\alpha(h)}v_{n}^{\beta(g)}u_{n}^{-\alpha(h)}v_{n}^{-\beta(g)}u_{n}^{-\alpha(g)}$
$=\omega^{-\beta(g)\alpha(h)}u_{n}^{\alpha(g)+\alpha(h)}u_{n}^{-\alpha(h)}v_{n}^{\beta(g)}v_{n}^{-\beta(g)}u_{n}^{-\alpha(g)}$
$=\omega^{-\beta(g)\alpha(h)}\operatorname{id}_{\mathbb{C}^{n}}.$
Lemma 3.2.
Let $\Gamma$ be a countable discrete group and let $[\sigma]\in H^{2}(\Gamma;\mathbb{Z})$ be a cohomology class with cocycle representative $\sigma$. Suppose that $\rho_{n}$ is a sequence of functions from $\Gamma$ to $k_{n}\times k_{n}$ unitaries so that
$\rho_{n}(gh)\rho_{n}(h)^{-1}\rho_{n}(g)^{-1}=\exp\left(\frac{2\pi i}{n}\sigma(g,h)\right)\operatorname{id}_{\mathbb{C}^{n}}.$
Then
$||\rho_{n}(g)\rho_{n}(h)-\rho_{n}(gh)||\leq 2\pi|\sigma(g,h)|/n;$
in particular, $\rho_{n}$ is asymptotically multiplicative in operator norm.
Now suppose that $[\sigma]$ pairs nontrivially with some homology class. Then $\rho_{n}$ is not perturbable to homomorphisms in operator norm.
Proof.
This follows the proof of [7, Theorem 3.20].
The first part follows from the fact that
$||\rho_{n}(gh)-\rho_{n}(g)\rho_{n}(h)||=||\rho_{n}(gh)\rho_{n}(h)^{-1}\rho_{n}(g)^{-1}-\operatorname{id}_{\mathbb{C}^{k_{n}}}||.$
Now we will show that for large enough $n$, $\rho_{n}$ is not close to any genuine representation of $\Gamma$ in operator norm on a particular finite subset of $\Gamma$. There is some 2-cycle $c\in C_{2}(\Gamma)$ written
$c=\sum_{i=1}^{N}x_{i}[a_{i}|b_{i}].$
so that
$\langle\sigma,c\rangle\neq 0.$
Then we compute that
$\langle\rho_{n},c\rangle$ $=\frac{1}{2\pi i}\sum_{j=1}^{N}x_{j}\operatorname{Tr}(\log(\rho_{n}(a_{j}b_{j})\rho_{n}(b_{j})^{-1}\rho_{n}(a_{j})^{-1}))$
$=\frac{1}{2\pi i}\sum_{j=1}^{N}x_{j}\operatorname{Tr}\left(\log\left(\exp\left(\frac{2\pi i}{n}\sigma(a_{j},b_{j})\right)\operatorname{id}_{\mathbb{C}^{k_{n}}}\right)\right)$
$=\frac{1}{2\pi i}\sum_{j=1}^{N}x_{j}\frac{2\pi i}{n}\sigma(a_{j},b_{j})\operatorname{Tr}(\operatorname{id}_{\mathbb{C}^{k_{n}}})$
$=\langle\sigma,c\rangle\frac{k_{n}}{n}$
$\neq 0.$
By Theorem 2.6 it follows that $\rho_{n}$ cannot be within $\frac{1}{24}$ of a genuine representation on the boundary set $\{a_{j},b_{j},a_{j}b_{j}\}_{j=1}^{N}$ and thus cannot be perturbed to a genuine representation.
Now we are ready to prove Theorem 1.1.
Proof.
Note that by the definition of the cup product $\sigma(g,h)=-\beta(g)\alpha(h)$ is a cocycle representative of the cohomology class $-\beta\smile\alpha=\alpha\smile\beta$.
First, we show asymptotic multiplicativity. In operator norm, this follows directly from Lemma 3.1 and Lemma 3.2. For the $p<\infty$ case, we use the same lemmas to show
$||\rho_{n}(gh)-\rho_{n}(g)\rho_{n}(h)||_{p}$ $\leq||\rho_{n}(gh)-\rho_{n}(g)\rho_{n}(h)||\cdot n^{1/p}$
$\leq 2\pi|\alpha(g)\beta(h)|\cdot n^{1/p-1}.$
This goes to zero for $p>1$.
Since $\alpha\smile\beta$ is non-torsion it must also pair nontrivially with a 2-homology class; to see this note that from the universal coefficient theorem [10, Theorem 53.1] we have a short exact sequence
[…]
Since $\Gamma$ is finitely generated $H_{1}(\Gamma;\mathbb{Z})\cong\Gamma/[\Gamma,\Gamma]$ [1, page 36] is finitely generated as well. Thus $\operatorname{Ext}(H_{1}(\Gamma;\mathbb{Z}),\mathbb{Z}))$ can be show to be torsion from [10, Theorem 52.3] and the table on [10] page 331.
Now it follows from Lemma 3.1 and Lemma 3.2 that $\rho_{n}$ is not perturbable in operator norm to a sequence of genuine representations. Since operator norm is smaller than any other unnormalized Schatten $p$-norm it follows that $\rho_{n}$ cannot be close to a genuine representation in any of these norms either. ∎
Remark 3.3.
Note that the finitely generated condition in Theorem 1.1 can be dropped if we require that $\alpha\smile\beta$ pairs nontrivially with a 2-homology class.
Acknowledgments
I would like to thank Eli Bashwinger for answering a question I had about the cohomology of Thompson and Thompson-like groups; his answer inspired me to make this note. I would also like to thank the Purdue math department for supporting me with the Ross-Lynn grant during the 2023-2024 academic year.